# Optimal Oblivious Load-Balancing for Sparse Traffic in Large-Scale Satellite Networks

### v1

@@ VI Numerical Results and Conclusion
In conclusion, we show the optimality of the LLB policy in load-balancing sparse traffic in a $N\times N$ torus. These results can also be straightforwardly extended to the general case of $N\times M$ tori with unequal capacities in the horizontal and vertical directions.
@@ Appendix A MaxLoad Equivalence of $\mathcal{D}_{k}$ and $\mathcal{D}^{\prime}_{k}$
as the optimal points of a linear objective over a convex polytope are at its vertices [bertsimas_linear_optimization]. To show that $\text{vertices}(\mathcal{D}_{k})\subset\mathcal{D}^{\prime}_{k}$, we invoke certain properties of vertices from [bertsimas_linear_optimization]. From [bertsimas_linear_optimization], any vertex (extreme point) $d^{*}$ of convex polytope $\mathcal{D}_{k}\subseteq\mathcal{R}^{N^{4}-N^{2}}$, $d^{*}$ cannot be written as a convex combination of two other points in $\mathcal{D}_{k}$. Now, consider $d\in\mathcal{D}_{k}$ and $d\notin\mathcal{D}^{\prime}_{k}$. We will show that it is possible to write $d$ as a convex combination of other points in $\mathcal{D}_{k}$. Since $d\notin\mathcal{D}^{\prime}_{k}$ but $d\in\mathcal{D}_{k}$, either the number of sources with non-zero traffic in $d$ is strictly greater than $k$, or the number of sinks with non-zero traffic in $d$ is strictly greater than $k$. Without loss of generality, assume that the number of sources is greater than $k$ (a similar line of reasoning will hold when we assume that number of sinks is greater than $k$ instead). Additionally, the total amount of traffic in the network must satisfy $\sum_{s,t}d_{s,s+t}\leq k$ and each source can produce at most 1 unit of traffic because $d\in\mathcal{D}_{k}$.
This tells us that there must exist at least two sources that produce less than 1 unit of traffic. This can be proven by contradiction. If there is at most one source that produces strictly less than 1 unit of traffic, then at least $k$ sources must source exactly 1 unit of traffic. Then, the total traffic in the network is strictly greater than $k$ units, which is a contradiction.
Now, let $S=\{s_{1},s_{2},...,s_{N_{s}}\}$ be the set of $N_{s}$ sources which totally contribute to at most 1 unit of traffic. As discussed, $\sum_{t}d_{s,s+t}<1$ for $s\in S$, and $\sum_{s\in S}\sum_{t}d_{s,s+t}\leq 1$.
We now construct a whole array of traffic matrices and show that $d$ can be expressed as a convex combination. Define $d^{(0)}$ as follows.
@@ Appendix C Reduced Version of the Oblivious Routing Program
Before we state the proof, we need to state a fairly trivial lemma, which will help in the proof.
Lemma 5.
Suppose $d\in\mathcal{D}_{k}$, and let $d^{\prime}$ be a traffic matrix such that $d^{\prime}_{s,s+t}\triangleq d_{\phi_{1}(s),\phi_{1}(s)+\phi_{2}(t)}$ for all $s,t$, where $\phi_{1}$ and $\phi_{2}$ are arbitrary automorphisms. Then, $d^{\prime}\in\mathcal{D}_{k}$.
It is easy to see that Lemma 5 is true. Since $\phi_{1}$ and $\phi_{2}$ are maps from $V\to V$, they just change the source and destination pairs of the original traffic matrix $d$ without altering the amount of traffic flow itself. Upon the action of automorphisms $\phi_{1}$ and $\phi_{2}$, it is easy to see that $d^{\prime}$ would still satisfy the inequality constraints of $\mathcal{D}_{k}$, since it is just a reordering of $d$.We are now ready for the proof of Lemma 3.
@@
Now, consider any $d^{(1)}\in\mathcal{D}_{k}$ and let us define $d^{(2)}$ be a traffic matrix such that $d^{(2)}_{s,s+t}=d^{(1)}_{-s,-s+t}$ for all $s,t$. From Lemma 5, $d^{(2)}\in\mathcal{D}_{k}$. Consequently,
$\sum_{s,t}d^{(1)}_{s,s+t}g^{t}_{s,s+e_{1}}$ $=\sum_{s,t}d^{(2)}_{-s,-s+t}g^{t}_{s,s+e_{1}}$
$=\sum_{s,t}d^{(2)}_{s,s+t}g^{t}_{-s,-s+e_{1}}$ (Reindexing nodes)
@@
Since this is true for any arbitrary $d^{(1)}\in\mathcal{D}_{k}$, we get $\max_{d\in\mathcal{D}_{k}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}\leq\theta$.
[…]
$\sum_{s,t}d^{(1)}_{s,s+t}f^{s,s+t}_{i,i+e}$ $=\sum_{s,t}d^{(1)}_{s,s+t}g^{\phi(t)}_{\phi(i-s),\phi(i-s)+e_{1}}$
$=\sum_{s,t}d^{(1)}_{i-\phi(s),i-\phi(s)+\phi(t)}g^{t}_{s,s+e_{1}},$
where the last equality comes from reindexing the summation. Let us define $d^{(2)}$ as $d^{(2)}_{s,s+t}=d^{(1)}_{i-\phi(s),i-\phi(s)+\phi(t)}$. From Lemma 5, $d^{(2)}\in\mathcal{D}_{k}$. Therefore,
$\sum_{s,t}d^{(1)}_{s,s+t}f^{s,s+t}_{i,i+e}$ $=\sum_{s,t}d^{(1)}_{i-\phi(s),i-\phi(s)+\phi(t)}g^{t}_{s,s+e_{1}}$
$=\sum_{s,t}d^{(2)}_{s,s+t}g^{t}_{s,s+e_{1}}$
@@ Appendix D Proof of Worst-Case Load Lower Bound
To prove part (a) of Theorem 1, we first consider the dual of the inner minimization problem $\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}$ for any traffic matrix $d\in\mathcal{D}_{k}$. We focus on the case where $N$ is even for simplicity. The case when $N$ is odd is similar, with minor modifications. From strong duality, the problem $\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}$ is equivalent to
$\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}=\max_{\nu}$ $\quad\frac{1}{4}\sum_{t}\nu^{t}_{t}$ (17)
s.t. $\quad\nu^{t}_{i+e}-\nu^{t}_{i}\leq c(d,t,i,e)$ $\forall i,t$
@@
where $\lambda(d,i,t,e)=d_{\phi(i),\phi(i)+\phi(t)}$ such that $e=\phi(e_{1})$ and $\phi\in\{I,R_{xy},R_{0},R_{0}R_{xy}\}$. Then, the problem $\max_{d\in\mathcal{D}_{k}}\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}$ is equivalent to
@@
This is very similar to the techniques used in [applegate_making_2006]. Refer to [applegate_making_2006] for more details on how to formulate the equivalent maximization problem. In problem (18), we interpret $c(d,t,i,e)$ as the cost of using edge $(i,i+e)$ in a route to $t$ under traffic matrix $d$. Due to the telescopic nature of constraint (18), we can conclude that for any path from 0 to $t$ (denoted as $\mathcal{P}_{t}$),
$\nu^{t}_{t}\leq\sum_{(i,i+e)\in\mathcal{P}_{t}}c(d,t,i,e)$
This has to be true for the minimum cost path as well, i.e., $\nu^{t}_{t}\leq\min_{\mathcal{P}_{t}}\sum_{(i,i+e)\in\mathcal{P}_{t}}c(d,t,i,e)$. Since we need to maximize the sum of $\nu^{t}_{t}$, we can greedily set $\nu^{t}_{t}=\min_{\mathcal{P}_{t}}\sum_{(i,i+e)\in\mathcal{P}}c(d,t,i,e)$, which consequently leads to
$\theta^{*}=\frac{1}{4}\max_{d\in\mathcal{D}_{k}}\min_{\bar{\mathcal{P}}}\sum_{t}\sum_{(i,i+e)\in\mathcal{P}_{t}}c(d,t,i,e),$
@@
Based on this observation, if we select a traffic matrix $d\in\mathcal{D}_{k}$ such that the cost of using edge $(i,i+e)$ is constant, i.e., $c(d,i,t,e)=c$ for a particular choice of $t$ and $\forall i\in B_{r}(0),\forall e$ directed away from 0 and $\forall i\in B_{r}(t),\forall e$ directed toward $t$, then we see that any path from 0 to $t$ would incur a cost of at least $2rc$.
The split diamond traffic matrix of size $r$ does exactly this. It selects $t$ to be the farthest node from the origin, which is $(N/2,N/2)$. The sources and destinations of traffic are carefully selected such that every hop that is directed away from the origin in $B_{r}(0)$ incurs a cost of $1$ and every hop that is directed toward $t$ in $B_{r}(t)$ incurs a cost of $1$. Hence, $c=1$ and every path from 0 to $t=(N/2,N/2)$ incurs a cost of $2r$. As mentioned earlier, for the split diamond traffic of size $r$, $\sum_{s,t}d^{SD(r)}_{s,s+t}=2r^{2}$ must be lesser than or equal to $k$ for $d^{SD(r)}$ to belong to $\mathcal{D}_{k}$. The sum cost of the objective under Split-Diamond traffic of size $r$ is $\frac{1}{4}\times 2r=\frac{r}{2}$. We maximize this sum cost under the constraint $2r^{2}\leq k,r\in\mathbb{N}$. When $2k$ is a perfect square, the best choice of $r$ would be $\sqrt{k/2}$ and the value of the objective is $\sqrt{2k}/4$. Since $d^{SD(r)}$ is a feasible solution, we arrive at $\theta^{*}\geq\sqrt{2k}/4$ when $2k$ is a perfect square.
@@
$\theta^{*}$ $=\max_{d\in\mathcal{D}_{k}}\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}}$
$\geq\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}\bar{d}_{s,s+t}g^{t}_{-s,-s+e_{1}}$
$\geq(1-\alpha)\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d^{SD(m)}_{s,s+t}g^{t}_{-s,-s+e_{1}}$
$\quad+\alpha\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d^{SD(m+1)}_{s,s+t}g^{t}_{-s,-s+e_{1}}$
@@ Appendix E Optimality of Local Load-Balancing Policy
$2r\times\frac{1}{4r}\times\sum_{i=0}^{r-1}(r-i)+(k-2r)\times\frac{1}{8r}=\frac{r}{4}+\frac{k}{8r}.$
@@
Appendix F Proof of Lemma 4
@@
The crux of the proof lies in invoking Menger’s Theorem [lawler_combinatorial_1976], which is a special case of the more general and celebrated Max-Flow Min-Cut Theorem. Before we prove the lemma, we define some preliminaries. For two mutually exclusive subsets of vertices, $S\subset V$ and $T\subset V$, we define the cut set (or simply, a cut) $C(S,T)$ as the set of edges that are directed from $S$ to $T$ such that, once removed, there exists no directed path from any node in $S$ to any node in $T$. We define the min-cut of $(S,T)$ as
@@
Appendix G Optimality of Valiant Load-Balancing for $k\geq N^{2}/2$

### v2

@@ VIII Numerical Results and Conclusion
In conclusion, we show the optimality of the LLB policy in load-balancing sparse traffic in a $N\times N$ torus.
@@ Appendix A MaxLoad Equivalence of $\mathcal{D}_{k}$ and $\mathcal{D}^{\prime}_{k}$
as the optimal points of a linear objective over a convex polytope are at its vertices [2]. To show that $\text{vertices}(\mathcal{D}_{k})\subset\mathcal{D}^{\prime}_{k}$, we invoke certain properties of vertices from [2]. From [2], any vertex (extreme point) $d^{*}$ of convex polytope $\mathcal{D}_{k}$ cannot be written as a convex combination of other points in $\mathcal{D}_{k}$. Now, consider a traffic matrix $d$ such that $d\in\mathcal{D}_{k}$ and $d\notin\mathcal{D}^{\prime}_{k}$. We will show that it is possible to write $d$ as a convex combination of other points in $\mathcal{D}_{k}$. Since $d\notin\mathcal{D}^{\prime}_{k}$ but $d\in\mathcal{D}_{k}$, either the number of sources with non-zero traffic in $d$ is strictly greater than $k$, or the number of sinks with non-zero traffic in $d$ is strictly greater than $k$. Without loss of generality, assume that the number of sources is greater than $k$ (a similar line of reasoning will hold when we assume that number of sinks is greater than $k$ instead). As $d\in\mathcal{D}_{k}$, the total amount of traffic in the network must satisfy $\sum_{s,t}d_{s,s+t}\leq k$ and each source can produce at most 1 unit of traffic because $d\in\mathcal{D}_{k}$.
This tells us that there must exist $N_{s}\geq 2$ sources that collectively produce less than 1 unit of traffic. This can be proven by contradiction. Assume there is at most $N_{s}=1$ source that produces strictly less than 1 unit of traffic. The other sources produce exactly 1 unit of traffic. Since there are at least $k$ other such sources, the total traffic in the network exceeds $k$ units, which is a contradiction.
Let $S=\{s_{1},s_{2},...,s_{N_{s}}\}$ be the set of $N_{s}$ sources which totally contribute to at most 1 unit of traffic. As a result of the previous paragraph, $\sum_{t}d_{s,s+t}<1$ for $s\in S$, and $\sum_{s\in S}\sum_{t}d_{s,s+t}\leq 1$.
We now construct a whole bunch of traffic matrices and show that $d$ can be expressed as a convex combination of these traffic matrices. Define $d^{(0)}$ as follows.
@@ Appendix C Reduced Version of the Oblivious Routing Program
$\sum_{s,t}d^{(1)}_{s,s+t}f^{s,s+t}_{i,i+e}$ $=\sum_{s,t}d^{(1)}_{s,s+t}g^{t}_{i-s,i-s+e}$
$=\sum_{s,t}d^{(1)}_{s+i,s+i+t}g^{t}_{-s,-s+e},$
where the last equality comes from reindexing the summation. Let us define $d^{(2)}$ as $d^{(2)}_{s,s+t}=d^{(1)}_{s+i,s+i+t}$. Note that $d^{(2)}\in\mathcal{D}_{k}$, as the $d^{(2)}$ is just the translation of source-destination pairs of $d^{(1)}$. Additionally, note that the set of directions $A=\{e_{1},e_{2},-e_{1},-e_{2}\}$ can be completely produced using $e_{1}$ and the $I,R_{xy},R_{0}$ and $R_{xy}R_{0}$ automorphisms respectively. Therefore, from automorphism invariance, we get
$g^{t}_{-s,-s+e}=g^{\phi(t)}_{\phi(-s),\phi(-s)+e_{1}}$,
where $\phi$ is the automorphism which maps $e$ to $e_{1}$, and,
$\sum_{s,t}d^{(1)}_{s,s+t}f^{s,s+t}_{i,i+e}$ $=\sum_{s,t}d^{(1)}_{s+i,s+i+t}g^{t}_{-s,-s+e}$
$=\sum_{s,t}d^{(2)}_{s,s+t}g^{t}_{-s,-s+e}$
$=\sum_{s,t}d^{(2)}_{s,s+t}g^{\phi(t)}_{\phi(-s),\phi(-s)+e_{1}}$
$=\sum_{s,t}d^{(2)}_{\phi(s),\phi(s)+\phi(t)}g^{t}_{-s,-s+e_{1}}.$
Again, the last equality follows from a reindexing of the summation under the automorphism $\phi$. If we define $d^{(3)}$ to be a traffic matrix such that $d^{(3)}_{s,s+t}=d^{(2)}_{\phi(s),\phi(s)+\phi(t)},$ where $\phi\in\{I,R_{xy},R_{0},R_{xy}R_{0}\}$, then observe that $d^{(3)}\in\mathcal{D}_{k}$ also. Finally,
$\sum_{s,t}d^{(1)}_{s,s+t}f^{s,s+t}_{i,i+e}$ $=\sum_{s,t}d^{(2)}_{\phi(s),\phi(s)+\phi(t)}g^{t}_{-s,-s+e_{1}}$
$=\sum_{s,t}d^{(3)}_{s,s+t}g^{t}_{-s,-s+e_{1}}$
@@ Appendix D Proof of Worst-Case Load Lower Bound
To prove part (a) of Theorem 1, we first consider the dual of the inner minimization problem $\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}}$ for any traffic matrix $d\in\mathcal{D}_{k}$. We focus on the case where $N$ is even for simplicity. The case when $N$ is odd is similar, with minor modifications. From strong duality, the problem $\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}$ is equivalent to
$\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}}=\max_{\nu}$ $\quad\frac{1}{4}\sum_{t}\nu^{t}_{t}$ (20)
s.t. $\quad\nu^{t}_{i+e}-\nu^{t}_{i}\leq\lambda(d,t,i,e)$ $\forall i,t$
@@
where $\lambda(d,i,t,e)=d_{-\phi(i),-\phi(i)+\phi(t)}$ such that $e=\phi(e_{1})$ and $\phi\in\{I,R_{xy},R_{0},R_{0}R_{xy}\}$. Then, the problem $\max_{d\in\mathcal{D}_{k}}\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}$ is equivalent to
@@
This is very similar to the techniques used in [1]. Refer to [1] for more details on how to formulate the equivalent maximization problem. In problem (21), we interpret $\lambda(d,t,i,e)$ as the cost of using edge $(i,i+e)$ in a route to $t$ under traffic matrix $d$. Due to the telescopic nature of constraint (21), we can conclude that for any path from 0 to $t$ (denoted as $\mathcal{P}_{t}$),
$\nu^{t}_{t}\leq\sum_{(i,i+e)\in\mathcal{P}_{t}}\lambda(d,t,i,e)$
This has to be true for the minimum cost path as well, i.e., $\nu^{t}_{t}\leq\min_{\mathcal{P}_{t}}\sum_{(i,i+e)\in\mathcal{P}_{t}}\lambda(d,t,i,e)$. Since we need to maximize the sum of $\nu^{t}_{t}$, we can greedily set $\nu^{t}_{t}=\min_{\mathcal{P}_{t}}\sum_{(i,i+e)\in\mathcal{P}}\lambda(d,t,i,e)$, which consequently leads to
@@
Based on this observation, if we select a traffic matrix $d\in\mathcal{D}_{k}$ such that the cost of using edge $(i,i+e)$ is constant, i.e., $\lambda(d,i,t,e)=\lambda$ for a particular choice of $t$ and $\forall i\in B_{r}(0),\forall e$ directed away from 0 and $\forall i\in B_{r}(t),\forall e$ directed toward $t$, then we see that any path from 0 to $t$ would incur a cost of at least $2r\lambda$.
The split diamond traffic matrix of size $r$ does exactly this. It selects $t$ to be the farthest node from the origin, which is $(N/2,N/2)$, only when $k\leq N^{2}/2$. The sources and destinations of traffic are carefully selected such that every hop that is directed away from the origin in $B_{r}(0)$ incurs a cost of $1$ and every hop that is directed toward $t$ in $B_{r}(t)$ incurs a cost of $1$. Hence, $\lambda=1$ and every path from 0 to $t=(N/2,N/2)$ incurs a cost of $2r$. As mentioned earlier, for the split diamond traffic of size $r$, $\sum_{s,t}d^{SD(r)}_{s,s+t}=2r^{2}$ must be lesser than or equal to $k$ for $d^{SD(r)}$ to belong to $\mathcal{D}_{k}$. The sum cost of the objective under Split-Diamond traffic of size $r$ is $\frac{1}{4}\times 2r=\frac{r}{2}$. We maximize this sum cost under the constraint $2r^{2}\leq k,r\in\mathbb{N}$. When $2k$ is a perfect square, the best choice of $r$ would be $\sqrt{k/2}$ and the value of the objective is $\sqrt{2k}/4$. Since $d^{SD(r)}$ is a feasible solution, we arrive at $\theta^{*}\geq\sqrt{2k}/4$ when $2k$ is a perfect square.
@@
$\theta^{*}$ $=\max_{d\in\mathcal{D}_{k}}\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}$
$\geq\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}\bar{d}_{s,s+t}g^{t}_{s,s+e_{1}}$
$\geq(1-\alpha)\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d^{SD(m)}_{s,s+t}g^{t}_{s,s+e_{1}}$
$\quad+\alpha\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d^{SD(m+1)}_{s,s+t}g^{t}_{s,s+e_{1}}$
@@ Appendix E Optimality of Local Load-Balancing Policy
$2\times\frac{1}{4r}\times\sum_{i=0}^{r-1}(r-i)+(k-2r)\times\frac{1}{8r}=\frac{r}{4}+\frac{k}{8r}.$
@@
Appendix F Modifications to LLB policy under specific conditions
As discussed in Section V, it suffices to describe the routing policy when $t_{x}\leq t_{y}\leq N/2$ as for other destinations, the routing can be obtained from automorphism transformations. In this section, we limit ourselves to $t_{x}\leq t_{y}\leq N/2$ for ease of presentation. (For modifications to the GLLB algorithm, just replace $S_{r}(0)$ and $S_{r}(t)$ with $S_{r_{1},r_{2}}(0)$ and $S_{r_{1},r_{2}}(t)$ respectively. The following arguments are valid for the GLLB routing algorithm as well).
F-1 Type-I Overlap
F-2 Type-II Overlap
Appendix G Proof of Lemma 4
@@
The crux of the proof lies in invoking Menger’s Theorem [11], which is a special case of the more general and celebrated Max-Flow Min-Cut Theorem. Before we prove the lemma, we define some preliminaries. For two mutually exclusive subsets of vertices, $S\subset V$ and $T\subset V$, we define the cut set (or simply, a cut) $C(S,T)$ as the set of edges that are directed from $S$ to $T$ such that, once removed, there exists no directed path from any node in $S$ to any node in $T$. We define the min-cut of $(S,T)$ as
@@
Appendix H Optimality of Valiant Load-Balancing for $k\geq N^{2}/2$